OpenStudy (zale101):

Find the magnitude of the resultant force and the angle that makes with the positive x-axis

OpenStudy (zale101):



OpenStudy (zale101):

\(F1=75 N; θ=0\) \(F2=50 N; θ=30\) \(F3=40 N; θ=150\) F1= \(f_1x=75cos(0)=75 N\) \(f_1y=75sin(0)=0 N\) F2= \(f_2x=50cos(30)=25\sqrt{3} N\) \(f_2y=50sin(30)=25 N\) F3= \(f_3x=40cos(150)=−20\sqrt{2}N\) \(f_3y=40sin(150)=20 N\)

OpenStudy (zale101):

Is this good so far?

jimthompson5910 (jim_thompson5910):

I don't agree with f3x

jimthompson5910 (jim_thompson5910):

everything else looks good

ganeshie8 (ganeshie8):

Looks good! next simply add them component by component

ganeshie8 (ganeshie8):

Ahh I didn't check the arithmetic..

ganeshie8 (ganeshie8):

this should fix it : \(F1=75 N; θ=0\) \(F2=50 N; θ=30\) \(F3=40 N; θ=150\) F1= \(f_1x=75cos(0)=75 N\) \(f_1y=75sin(0)=0 N\) F2= \(f_2x=50cos(30)=25\sqrt{3} N\) \(f_2y=50sin(30)=25 N\) F3= \(f_3x=40cos(150)=−20\sqrt{\color{red}{3}}N\) \(f_3y=40sin(150)=20 N\)

jimthompson5910 (jim_thompson5910):

much better

OpenStudy (zale101):

Okay. Thanks! @jim_thompson5910 and @ganeshie8 !!! I'll deal with the arithmetic, it's not the problem for me now. But the next step is what i'm having problems with.

jimthompson5910 (jim_thompson5910):

you have the vector components, add up the corresponding components

jimthompson5910 (jim_thompson5910):

u = <a,b> v = <c,d> w = <e,f> u+v+w = <a+c+e, b+d+f>

ganeshie8 (ganeshie8):